OpenStudy (anonymous):

To multiply or divide a number by 10,000, how many places should you move the decimal point? four three two one

OpenStudy (freckles):

say we the number 5 and we multiply it by 10 we get 50 (decimal places given one more) and we multiply it by 100 we get 500 (decimal places given two more) and we multiply it by 1000 we get 5000 (decimal places given 3 more)

OpenStudy (freckles):

so what do you think

OpenStudy (chosenmatt):

yeah what do you think?

OpenStudy (jhannybean):

When dividing by 10000, count the number of 0's there are and move the decimal place to the left. When multiplying a number by 10000, move the decimal place to the right the same number of 0's given.

OpenStudy (zzr0ck3r):

yet another way of looking at it 5/1 = 5.0 5/10 = 0.5 what do you think 5/100 = ?

OpenStudy (jhannybean):

double posting to show more emphasis? :P

OpenStudy (zzr0ck3r):

I made it 5.0 instead of 5 I guess the answer is yes, because I did that to emphasis the moving of the decimal. ;)